O
Please answer steps by steps
Divine Bath Works is a leader in the bath fixtures industry, rising to prominence with the introduction of over-the-counter sinks. The following data pertain to September, the last month of Divines fiscal year. Type of Inventory Balance as of September 1 Materials (direct and indirect) Work in Process Finished Goods $28,100 $124,320 $245,000 The Work in Process Inventory on 9/1 consisted of one job, Job X, for 700 units of Model OO, an oval, over-the-counter sink. The Finished Goods Inventory on 9/1 comprised the following: Model Number (description) O0 (Oval, over-the counter) RI (Round, in-the counter) OI (Oval, in-the counter) Quantity Cost per unit Total Cost 100 400 450 $320.00 $197.25 $298.00 $32,000 $78,900 $134,100 $245,000 Total Divine uses a normal costing system, combined with the FIFO inventory flow assumption, to cost its inventories. That is, Divine traces actual materials and labor costs to individual jobs. It allocates manufacturing overhead using a predetermined overhead rate computed at the start of the year. Divine allocates overhead using direct labor hours as the allocation basis (cost driver).
Divine prepared the following annual overhead cost budget at the start of the present year. Overhead item Supplies Indirect Labor Supervision Depreciation Utilities Budgeted amount $290,000 422,340 228,520 425,800 136,740 185,000 Factory Rent Total $1,688,400 Divine also informs you that it had budgeted for 80,400 direct labor hours at the start of the year. As of September 1, the company had incurred 75,315 actual direct labor hours and spent $1,569,450 in actual overhead costs. During September, Divine purchased $112,340 of direct materials and $26,430 of indirect materials. The following table shows the requisition of direct materials during September: lob number Model Description Units OI Oval, in sink 500 RO Oval, OTC 700 Round, OTC Round, in sink Direct Materials $22,000 $58,000 200 $24,000 700 $11,450 Divine consumed $22,000 of indirect materials during September. Divine incurred the following labor costs during September. Item Indirect Labor Direct Labor Amount $32,000 $34,560 $43,740 $29,365 $15,210 Job Hours 1,920 2,430 1,678 845
The following comprises actual overhead costs (other than supplies and indirect labor) incurred during the month of September. Cost Item Supervision Depreciation Utilities Factory rent Actual amount (for September) $18,400 $32,650 $ 8,900 $15,400 Finally, Divine completed all jobs except Job B during September. It also sold the following items during the month of September. Model Units Sold RI 200 OI 400 RO 150 750 REQUIRED: 1. Compute the amount of under-over applied overhead for the month of September and the year ended September 30. September: Year: (10 pts.) 2. Compute the ending balance of the Work in Process Inventory account at 9/30: (5 pts.) 3. Calculate the ending balance of Finished Goods Inventory for each model at 9/30 (use FIFO method). (20 pts.) Model 0O: Model RI: Model OI: Model RO:
4. Prepare an Income Statement for the month ended 9/30 (Assume that Sales Revenue will be 25% greater than total Cost of Goods Sold, operating expenses are $20,000 and the companys income tax rate is 20%). (10 pts.) 5. Team Participation Points: these will be awarded by your teammates (15 pts.)
Answer #1
1 Predetermined overhead rate=Estimated overhead cost/Estimated direct labor hours=1688400/80400=$21 per DLH
For the month of September:
Actual overhead cost:
$
Indirect materials 22000
Indirect labor 32000
Supervision 18400
Depreciation 32650
Utilities 8900
Factory rent 15400
Total 129350
Applied overhead cost=Actual direct labor hours* Predetermined overhead rate
Actual direct labor hours:
Hours
Job X 1920
Job K 2430
Job L 1678
Job B 845
Total 6873
Applied overhead cost=6873*21=$ 144333
Applied overhead > Actual overhead. Hence over application of overhead
Over applied overhead=144333-129350=$ 14983
For the year ended September 30:
Actual overhead cost=Overhead incurred till Sep 1+ Overhead incurred durig september=1569450+129350=$ 1698800
Applied overhead cost=Actual direct labor hours* Predetermined overhead rate
Actual direct labor hours=DLH till Sep 1+ DLH durig september=75315+6873=82188 hours
Applied overhead cost=82188*21=$ 1725948
Applied overhead > Actual overhead. Hence over application of overhead
Over-applied overhead=1725948-1698800=$ 27148
2 Ending balance of work in process cosnsist of Cost of jobs not completed
Here, the only job which is not completed is Job B
Cost of Job B will give the ending balance in work in process
Cost of Job B:
$
Direct materials 11450
Direct labor 15210
Overhead applied (845*21) 17745
Total job cost 44405
Ending balance of work in process=$ 44405